Mars Society Flashline Arctic Research Station Mission Support
Home About Team Mission Logs FMARS Home RMMS Home

Journalist's Report for July 19, 2001, 7:15pm MDT

So far we've not received any reports from the Field Team. Our hope is that they have been busy installing and repairing the Coms equipment and opening packages which were scheduled for landing several days ago.

We did send the Space Weather Report for today and a very brief News From Earth Selection. Lorraine Bell, Mission Director and Acting Journalist decided against sending news reports about various battles, deaths, and maulings so the stories to choose from were pretty skimpy.

Yesterday's reports indicated that the field team still had not found the ice axe holsters that had been included in the space suit packages. We sent a more detailed message seeking answers about the number of boxes received at the Hab and describing how the axe holsters were packed. Lorraine has a copy of the contents of all the shipping containers but unfortunately it is not filed here at Mission Support so couldn't be accessed. (DOH!)

At approximately 3:30pm MDT, we received a 'space line' call from Dr. Zubrin. The field team was concerned that messages from the KSC "public question and answer session" had not been received. It turned out that KSC had a terrible storm and the public decided to evacuate. We look forward to a sucessful session tomorrow.

At 7:30pm MDT we received an audio status report from the Field Team. They reported that all the crew was well and that the Hab was in generally good shape. However, the power problems continue. The hab needs 10kW of power and is currently using two generators but they continually lose power which causes problems with their electronic equipment - much of which is very expensive and delicate. We are looking into purchasing a 10kW diesel generator which could be sent to them.

The team also reported on their 3 person EVA. Zubrin, Frankel, and Bos all went on an EVA to explore the area to the East along the crater rim. Cathrine Frandsen was their CapCom inside the hab. The suiting up process is getting easier for the crew, and thus took less time. They expect to have it down pat pretty soon.

They reported that they continued to have some radio problems - probably the previous cold weather affecting the batteries - but they were able to successfully complete a 2-hour EVA. Cathrine recommended the use of personal digital recorders by all the EVA team. That would help document the EVA even if there was a radio failure. She also noted that the CapCom could use a map so that they could follow the EVA. We certainly concur with this as we have been frustrated by our own lack of a map!

While on the EVA, the team was treated to magnificent views - yes we're jealous. They also collected some very interesting samples: in an area of sedimentary matrials they found what looks to be some breccia (a sample of which they will cut for further analysis), in Lowell canal they took water and snow samples, and Brent Bos found a piece of limestone that may have one been a part of the coral reef.

There current plan is to head North of the Hab tomorrow, crossing Von Braun Planitia. On the 21st, they hope to head West. The goal is to create an operational map of the area around the Hab and then to choose interesting sites worthy of further study.

AS of 8:00pm MDT we had not received any further reports and so we signed off with the field team.

Lorraine Bell, Mission Support Director and Journalist

Requests concerning this web site should be directed to the Webmaster.
Mars Illustration Copyright 1999 by Mike Koonce/Mars Society
Copyright © 2001 The Mars Society. All rights reserved.
Mars Society Flashline Arctic Research Station Mission Support